OpenStudy (anonymous):

The pyramids at Giza include several square-based pyramids that were built in honor of the Egyptian gods. One of these incredible monuments is the Khafre pyramid, built as part of a temple complex, with the famous Sphinx at its base. The base length of this pyramid is 705 feet, and its height is 471 feet. Massive dimensions indeed! Calculate the volume of the Khafre pyramid. Show the steps of your solution and explain your work.

OpenStudy (ghazi):

area of base= (705)^2=497025 and area of pyramid = 497025*471*1/3=78032925

OpenStudy (anonymous):

thanks

OpenStudy (ghazi):

:)